OpenStudy (anonymous):

Functions f(x) and g(x) are described as follows: f(x) = -2x2 + 3 x g(x) 0 0 1 2 2 3 3 2 4 0 Which statement best compares the maximum value of the two functions?

OpenStudy (anonymous):

Any help would be much appreciated.

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Are busy?

Vocaloid (vocaloid):

list of statements?

OpenStudy (anonymous):

Oh... my bad. I knew I was forgetting something.

OpenStudy (anonymous):

Which statement best compares the maximum value of the two functions? It is equal for both functions. It is 2 units higher for f(x) than g(x). It is 2 units lower for f(x) than g(x). It is 1 unit higher for f(x) than g(x).

Vocaloid (vocaloid):

ok, so let's just look at f(x) for now f(x) = -2x^2 + 3 what is the maximum value of f(x)?

OpenStudy (anonymous):

Okay.. sorry for the long response. Open Study isn't agreeing with me... as usual.

OpenStudy (anonymous):

aah geez. I can't remember how to do f(x) problems...

Vocaloid (vocaloid):

hint: the maximum value of f(x) happens when x = 0 so, when x = 0, what is f(x)?

OpenStudy (anonymous):

Not there. I thought x = 1. So if x = o than... f(x) would just be... 0... I am sure I'm wrong about that lol. I have never been good at Algebra.

Vocaloid (vocaloid):

f(x) = -2x^2 + 3 = -2(0) + 3 = ?

OpenStudy (anonymous):

uh...

Vocaloid (vocaloid):

-2(0) + 3 = ?

OpenStudy (anonymous):

3?

Vocaloid (vocaloid):

right, so our maximum value of f(x) is 3 now take a look at g(x), what's the maximum value?

OpenStudy (anonymous):

Whats the equation? Is it still: f(x) = -2x2 + 3 only with g(x) in place of f(x)

Vocaloid (vocaloid):

no, it's the table...

Vocaloid (vocaloid):

it's this part x g(x) 0 0 1 2 2 3 3 2 4 0

OpenStudy (anonymous):

The highest number on the table is 3.

Vocaloid (vocaloid):

right, so the maximum value of g(x) is 3 so our final answer is...?

OpenStudy (anonymous):

uh...

OpenStudy (anonymous):

D!

OpenStudy (anonymous):

no... its A. lol

Vocaloid (vocaloid):

right, A

OpenStudy (anonymous):

A medal for you and a....medal for me...?

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

Yay! Now I have a wopping 3 medals lol. Thanks for all your help.